What is the answer to
3y^(-1/2) = 6

Answer by checkley77(12569)

3y^(-1/2) = 6
3/sqrty=6 square both sides.
9/y=36 cross multiply
36y=9
y=9/36
y=1/4 ans.
Proof:
3*1/4^(-1/2)=6
3/sqrt1/4=6
3/(1/2)=6
3*2/1=6
6=6